OpenStudy (anonymous):

I will fan and medal if you help me!!!!!! There is 5 quetions and if you help you will get a fan and medal Find the interquartile range of the data. A. 49 B. 35 C. 16 D. 14

OpenStudy (anonymous):

Find the median of the lower and upper half of your data. The median is the "midpoint," or the number that is halfway into a set. ... Subtract Q3 - Q1 to determine the IQR. Now you know how many numbers lie between the 25th percentile and the 75th percentile.

OpenStudy (anonymous):

31,33,35,41,43,46,48,49,49,50 Q2 (median) = (43 + 46)/2 = 89/2 = 44.5 Q1 = 35 Q3 = 49 Interquartile range is Q3 - Q1 = 49 - 35 = 14